Diamond is a natural form of pure carbon .What number of atoms of carbon are in a 1.0 carad diamond


1
Expert's answer
2021-08-31T01:32:17-0400

1.00 carat = 0.200 g

1 mol contains 6.022×1023 elementary entities of the substance.

n = m / M

M (C) = 12.0107 g/mol

n (C) = 0.200 / 12.0107 = 0.017 moles

Number of atoms = 0.017 x 6.022×1023 = 0.102 x 1023